problem 1) Ninety students will graduate from Lima Shawnee High School this spring. Of the 90 students, 50 are planning to ttend college. Two students are to be picked at random to carry flags at graduation.
A) What is the probability both of the selected students plan to attend college?
B) What is the probability one of the two selected students plans to attend college?
I know that these are counting problems that involve the basic counting rule, permutations and combinations. I need the formulas and answers. I am extremly lost.
problem 2) a survey of undergraduate students in the Shool of Business at Northern University revealed the following regarding the gender and majors of the students.
Gender Accounting Management Finance total
Male 100 150 50 300
Female 100 50 50 200
total 200 200 100 500
A) What is the probibility of selecting a female student?
B) What is the probability of selecting a finance or accounting major?
C) What is the probability of selecting a female or an accounting major? Which rule of addition did you apply?
D) Are gender and major independent? Why?
E) What is the probability of selecting an account major, given that the person selected is a male?
F) Suppose two students are selected randomly to attend a lunch with the president of the university. What is the probability that both of those selected are accounting majors?
Problem 3) Reynolds Construction Company has agreed not to erect all "look-alike" homes in a new subdivision. Five exterior designs are offered to potential home buyers. The bulider has standardized three interior plans that can be incorporated in any of the five exteriors. How many different ways can the exterior and interior plans be offered to potewntial home buyers?
Problem 4) A puzzle in the newspaper presents a matching problem. The names of 10 in one U.S. presdents are listed in column, and their vice presidents are listed in random order in the second column. The puzzle asks the reader to match each president with his vice president. If you make the matches randomly, how many matches are possible? What is the probability all 10 of your matches are correct?
Thank you in advance for this much needed help.

Answer by stanbon(75887)

1) Ninety students will graduate from Lima Shawnee High School this spring. Of the 90 students, 50 are planning to ttend college. Two students are to be picked at random to carry flags at graduation.
A) What is the probability both of the selected students plan to attend college?
Ans: 50C2/90C2
---------------------------
B) What is the probability one of the two selected students plans to attend college?
P(one of two) = 1 - [P(both do)+P(both don't)]
Ans: 1 - [[50C2/90C2 + 40C2/90C2]]
===============================================
problem 2) a survey of undergraduate students in the Shool of Business at Northern University revealed the following regarding the gender and majors of the students.
Gender Accounting Management Finance total
Male 100 150 50 300
Female 100 50 50 200
total 200 200 100 500
A) What is the probibility of selecting a female student?::200/500
-----------------------
B) What is the probability of selecting a finance or accounting major?
C) What is the probability of selecting a female or an accounting major?
Ans: (200 + 200 - 100)/500 = 300/500
---
Which rule of addition did you apply?
P(A or B) = P(A)+P(B)- P(A and B)
-----------------------------------------------
D) Are gender and major independent? Why?
Comment: What method is available to you for
determining independence?
-----------------------------------------------
E) What is the probability of selecting an account major, given that the person selected is a male?
---
P(accout|male) = 100/300 = 1/3
-----------------------------------------
F) Suppose two students are selected randomly to attend a lunch with the president of the university. What is the probability that both of those selected are accounting majors?
Ans: 200C2/500C2
------------------------
Problem 3) Reynolds Construction Company has agreed not to erect all "look-alike" homes in a new subdivision. Five exterior designs are offered to potential home buyers. The bulider has standardized three interior plans that can be incorporated in any of the five exteriors. How many different ways can the exterior and interior plans be offered to potewntial home buyers?
---
Ans: 5*3 = 15 plans
==========================
Problem 4) A puzzle in the newspaper presents a matching problem. The names of 10 in one U.S. presdents are listed in column, and their vice presidents are listed in random order in the second column. The puzzle asks the reader to match each president with his vice president. If you make the matches randomly, how many matches are possible?::: 2^10 = 1024 are possible
--------------------------------------------------
What is the probability all 10 of your matches are correct?
Ans 1/1024
===============
Cheers,
Stan H.